OpenStudy (anonymous):

The angle of depression from D to F measures 40 degrees. If EF= 14 yd, find DE. Round your answer to the nearest tenth.

OpenStudy (anonymous):

please help!!!

OpenStudy (aonz):

is there a diagram?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

just a triangle that is right

OpenStudy (aonz):

 like this?

OpenStudy (aonz):

why does the diagram stuff up??? :P

OpenStudy (aonz):

tan40 = 14/x rearrange to x = 14/ tan40 x= 16.68

OpenStudy (anonymous):

yes, and thank you so much